You are going to drive from Blackwood to Boston,a total trip of 345 miles,stopping at a couple of cities along the way to visit friends.The shortest leg of the trip will be the first one from Blackwood to Trenton,while the drive from Trenton to New York city will be 25 miles longer.The Final leg of the trip from New York to Boston will be double the trip from Blackwood to New York.How many miles is it between New York and Boston?
Answer by ikleyn(52921)

.
x + (x+25) + 2*(x+25) = 345   (where x = the distance from Blackwood to Trenton)


4x + 75 = 345


4x = 345 - 75 = 260  ====>  x = 260/4 = 65


Answer.  The distance under the question is  2*(x+25) = 2*(65+25) = 180 miles.

Internet (https://www.distancefromto.net/distance-from-new-york-to-boston-us) gives 190 miles.